if this is december,then last month was november.if last month was was november,then six months ago it was june.if six months ago it was june,then eleven months ago it was january.if next month will be january,then this is december.last month was november.Therefore, this is december{D,N,J,A,X}.Sir, please guide how to check the validity of this argument.THIS PROBLEM is on my agenda since a few years. Please Help.
Answer by Fombitz(32388)

Instead of names, let's use numbers.
1-January
2-February
.
.
.
11-November
12-December
So, if M=12, then M-1=11. True
If M-1=11, then M-6=M-1-5=11-5=6. True
If M-6=6, then M+1=M-6+7=6+7=13=1. True (the months repeat after 12 months so 13=1)
If M+1=13, then M+1-1=13-1=12. True